OpenStudy (bgrg007):

Algorithms prove that log(n(factorial)) = theta of (nlogn) is true, using Stirling's approximation, which is n(factorial)=(sqrt(2pi*n)) * ((n/e)^n) * (1 + theta of (1/n))

OpenStudy (bgrg007):

algorithm question*

OpenStudy (blockcolder):

\[\log n!=\Theta(n\log{n})\] Is this what you mean?

OpenStudy (blockcolder):

And Stirling's approximation is \[n! \sim \sqrt{2\pi n} \left(\frac{n}{e}\right)^n \left(1+\Theta\left(\frac{1}{n}\right)\right)\]

OpenStudy (bgrg007):

yes